A firm has a production function y = X1X2. If the minimum cost of production at w1 = w2 = 1 is equal to 4, what is y equal to?
1
Expert's answer
2020-11-09T07:05:02-0500

As per the given question,

"y=x_1 x_2"

Now,

"L=w_1x_1+w_2x_2+\lambda(y-x_1x_2)"

"0=\frac{dy}{d\lambda}=y-x_1x_2"


"0=\frac{dy}{d\lambda}=w_1-\lambda x_2"


"0=\frac{dy}{d\lambda}=w_2-\lambda x_2"

So, "\lambda = \frac{w_1}{x_2}=\frac{w_2}{x_1}"

We know that "4=w_1x_1+w_2x_2"

"4=w_1x_1+\frac{w_1\times x_1}{w_2}w_2"

"x_1=\frac{4}{2w_1}=\frac{2}{w_1}"

"x_2=\frac{2}{w_2}"

Hence, "y=x_1 x_2"

"y=\frac{2\times 2}{w_1\times w_2}"

"y=4"